Suppose a sample of 1635 floppy disks is drawn. Of these disks, 228 were defective. Using the data, construct the 98% confidence interval for the population proportion of disks which are defective. Round your answers to three decimal places.
Answer #1

Please don't hesitate to give a "thumbs up" for the answer in case the answer has helped you

p^ = x/n = 228/1635 = .13945

98% CI is given by a Z of 2.326

So, the 90% CI is given by

= p +/- Z*sqrt(p*p'/n)

= .13945 +/- 2.326*sqrt(.13945*.82055/1635)

= 0.120 to 0.159